### v1

# Symmetrizable systems

Abstract
Transforming an asymmetric system into a symmetric system makes it possible to exploit the simplifying properties of symmetry in control problems. We define and characterize the family of symmetrizable systems. These systems can be transformed into symmetric systems by a linear transformation of their inputs and outputs. We show that a Khatri-Rao rank needs to be satisfied for a system to be symmetrizable. We conclude that linear systems are generically neither symmetric nor symmetrizable.
I Introduction
Symmetric systems, also known as reciprocal systems in the literature, are control systems that exhibit input-output symmetry in their transfer function matrices $G(s)\in\mathbb{C}^{m\times m}$ as follows
$\Sigma_{e}G^{\top}(s)=G(s)\Sigma_{e},$ (1)
where $\Sigma_{e}$ is a constant diagonal matrix whose diagonal elements are either $1$ or $-1$. Relation (1) implies that a system is symmetric, if and only if, either its transfer function is a symmetric matrix or it becomes one by multiplying one or more of its columns by $-1$.
Symmetric systems are found in diverse application areas, such as electrical circuits, chemical reactors, mechanical systems, and power networks [mohammadpour2010efficient, pates2022passive]. These systems have properties that simplify many control problems. For example, certain $H_{2}$ and $H_{\infty}$ optimal control problems have known analytic solutions when applied to symmetric systems [pates2022passive, pates2019optimal], the optimal linear-quadratic regulator of symmetric systems can be obtained by iterative learning control with no bias [taghavian2024optimal], and estimated by a single trajectory of the system [drummond2024learning].
In this paper, we extend the family of symmetric systems, by introducing symmetrizable systems. These systems may not be symmetric, but after a linear transformation of their inputs and outputs as
$H(s)=K^{-1}G(s)K,$ (2)
they become symmetric, where $K\in\mathbb{R}^{m\times m}$ is a constant matrix, called the symmetrizing gain. This definition extends the simplifying properties of symmetric systems to asymmetric systems. Namely, one can first transform the asymmetric system $G(s)$ to a symmetric system $H(s)$ as (2), solve a control problem for $H(s)$, and convert the result back for $G(s)$. For example, consider the static output-feedback controller $u_{s}(t)=K_{s}\,y_{s}(t)$ designed for the symmetric system $H(s)$ with input $u_{s}$ and output $y_{s}$. This control law is equivalent to applying the following static output feedback to $G(s)$:
$u(t)=KK_{s}K^{-1}y(t),$ (3)
where $u$ and $y$ are the input and output of the asymmetric system $G(s)$, respectively.
We study the transformation (2) and provide the necessary and sufficient conditions for a given system $G(s)$ to admit such a transformation. Further, we propose a method to find a symmetrizing gain $K$ when one exists. The preliminaries are given in Section II and a background on symmetric systems is provided in Section III. Symmetrizable systems are introduced in Section IV and further discussed in Section V. We give some examples to demonstrate the results in Section VI and provide the concluding remarks in Section VII.
II Notation and preliminaries
We denote the algebraic multiplicity of an eigenvalue $\lambda$ by $\textnormal{mul}(\lambda)$, interpret $|M|$ and $\textnormal{sgn}(M)$ element-wise, and use $\textnormal{ker}(M)$ and $\textnormal{col}(M)$ to denote the kernel and column spaces of $M$. The direct sum of matrices is denoted by $\oplus$ and the Kronecker product is denoted by $\otimes$. The (column-wise) Khatri-Rao product of two matrices $M$ and $N$ partitioned as
$M$ $=\begin{bmatrix}M_{1}&M_{2}&\dots&M_{r}\end{bmatrix},\quad M_{j}\in\mathbb{R}^{m\times t_{j}}$
$N$ $=\begin{bmatrix}N_{1}&N_{2}&\dots&N_{r}\end{bmatrix},\quad N_{j}\in\mathbb{R}^{n\times s_{j}}$
is defined as
$M\star N:=\begin{bmatrix}M_{1}\otimes N_{1}&M_{2}\otimes N_{2}&\dots&M_{r}\otimes N_{r}\end{bmatrix}.$
Let $n_{+}(S)$, $n_{-}(S)$, and $n_{0}(S)$ be the number of positive, negative, and zero eigenvalues of the symmetric matrix $S\in\mathbb{R}^{q\times q}$ respectively. Then the signature of $S$ is defined by the scalar
$i(S):=n_{+}(S)-n_{-}(S).$
A diagonal matrix $S\in\mathbb{R}^{q\times q}$ is called a signature matrix if its diagonal elements are either $-1$ or $+1$. The two symmetric matrices $S_{1},S_{2}\in\mathbb{R}^{q\times q}$ are congruent, denoted by $S_{1}\sim S_{2}$, if there is a non-singular matrix $M$ such that $S_{2}=M^{\top}S_{1}M$. The centralizer of the square matrix $M$, denoted by $\textnormal{Cent}(M)$, is the set of matrices that commute with $M$. The following proposition shows that the centralizer of the direct sum of two matrices $M_{1}$ and $M_{2}$ can be obtained by the direct sum of their centralizer elements.
Proposition 1 ([LMU])
Let $M_{1}$ and $M_{2}$ be two square matrices with disjoint spectra (they have no common eigenvalues). Then $\textnormal{Cent}(M_{1}\oplus M_{2})=\{U_{1}\oplus U_{2}|\,U_{1}\in\textnormal{Cent}(M_{1}),U_{2}\in\textnormal{Cent}(M_{2})\}$.
III Symmetric systems
A linear system with transfer function $G(s)\in\mathbb{C}^{m\times m}$ is said to be (externally) symmetric if there is a signature matrix $\Sigma_{e}$ such that (1) holds [willems1976sym]. The matrix $\Sigma_{e}$ is called the external signature matrix of the system. The state-space
$\begin{array}[]{rcl}\dot{x}(t)&=&Ax(t)+Bu(t)\\ y(t)&=&Cx(t)+Du(t)\end{array},$ $\quad t\geq 0$ (6)
is called (internally) symmetric, if its associated short-hand system matrix
$P:=\begin{bmatrix}A&B\\ C&D\end{bmatrix}$ (7)
satisfies $\Sigma P=P^{\top}\Sigma$ for some signature matrix
$\Sigma=\begin{bmatrix}-\Sigma_{i}&0\\ 0&\Sigma_{e}\end{bmatrix}.$ (8)
We assume the state-space (6) represented by $P$ in (7) is minimal throughout the paper. In (8), we call $\Sigma$ the signature matrix of the system or system inertia, $i(\Sigma)$ the system signature, and $\Sigma_{i}$ the internal signature matrix of the system. Symmetric systems with the external and internal signature matrices $\Sigma_{e}=I$ and $\Sigma_{i}=-I$ are called completely symmetric [willems1976sym]. Relaxation systems are a well-known subset of these systems.
It is straightforward to verify that a system described by a symmetric state space with inertia (8) is symmetric with the external signature matrix $\Sigma_{e}$. The following proposition shows the converse is also true.
Proposition 2
[willems1976sym] Every symmetric system admits a symmetric minimal state-space realization.
IV Symmetrizable systems
In this section, we introduce a new family of linear systems, which we call symmetrizable systems. A linear system with transfer function $G(s)\in\mathbb{C}^{m\times m}$ is called symmetrizable if there exists a static gain $K\in\mathbb{R}^{m\times m}$ such that the transformed system (2) is symmetric. The transformation (2) induces the following equivalence relation which classifies all $m\times m$ transfer functions: $G\stackrel{{\scriptstyle\mathclap{\footnotesize\mbox{s}}}}{{\sim}}H$ is true if there is some non-singular $K\in\mathbb{R}^{m\times m}$ such that (2) holds. Hence, if a member of a class is symmetric, all members of that class are symmetrizable. If no member of a class is symmetric, then none of the members are symmetrizable.
Symmetrizable systems include symmetric systems as a subset. This follows from choosing $K=I$ in (2). The following theorem provides the conditions that characterize all symmetrizable systems.
Theorem 1
The system represented by (7) is symmetrizable if and only if the following equations have a non-singular solution for $Q$:
$PQ$ $=QP^{\top},\,Q=Q^{\top}$ (9)
$Q_{12}$ $=0.$ (10)
[…]
Proof:
Let $P$ in (7) be a system matrix of $G(s)$. Then all the minimal realizations of $H(s)=K^{-1}G(s)K$ are represented by the system matrix
$\begin{bmatrix}T^{-1}AT&T^{-1}BK\\ K^{-1}CT&K^{-1}DK\end{bmatrix}=L^{-1}PL,$
where
$L:=\begin{bmatrix}T&0\\ 0&K\end{bmatrix}$ (11)
is non-singular. Therefore, according to Proposition 2, system $G(s)$ is symmetrizable if and only if there is some block-diagonal matrix $L$ as (11) and some signature matrix $\Sigma$ such that
$\Sigma L^{-1}PL=L^{\top}P^{\top}L^{-\top}\Sigma.$ (12)
Define $Q:=L\Sigma L^{\top}$, which we partition as
$Q=\begin{bmatrix}Q_{11}&Q_{12}\\ Q_{12}^{\top}&Q_{22}\end{bmatrix},$
where $Q_{11}\in\mathbb{R}^{n\times n}$, $Q_{12}\in\mathbb{R}^{n\times m}$, $Q_{22}\in\mathbb{R}^{m\times m}$. Equation (12) can be written as (9)-(10) where
$Q_{11}$ $=-T\Sigma_{i}T^{\top},$ (13)
$Q_{22}$ $=K\Sigma_{e}K^{\top}.$ (14)
System $G(s)$ is symmetrizable with the internal and external signature matrices $\Sigma_{i}$ and $\Sigma_{e}$, if and only if (9), (10), (13), (14) have non-singular solutions for $Q$, $L$, and $K$. However, if a non-singular solution $Q$ exists for the linear equations (9) and (10), one can always find some $K$ and $T$ satisfying (13) and (14). To see this, note that since $Q$ is block-diagonal and non-singular, the submatrices $Q_{11}$, $Q_{22}$ are also non-singular. Therefore, there are signature matrices $\Sigma_{i}$, $\Sigma_{e}$ such that
$Q_{11}\sim-\Sigma_{i}\textnormal{ and }Q_{22}\sim\Sigma_{e}.$ (15)
By Sylvester’s law of inertia, equations (13) and (14) have solutions for $T$ and $K$ if and only if $\Sigma_{i}$, $\Sigma_{e}$ are chosen as (15). ∎
Remark 1
By setting an identity gain $K=I$, the equations (9), (10), (13), (14) provide the necessary and sufficient conditions for $G$ to be symmetric and characterize all its minimal symmetric realizations. These conditions are equivalent to the reciprocity conditions provided in [balancing].
When the system represented by (7) is symmetrizable, one can find a symmetrizing gain $K$ and a symmetric realization ($T$) based on the solution $Q$ of the equations (9)-(10). Such matrices $T$ and $K$ are not unique, with one instance being
$T=F_{1}|D_{1}|^{1/2},\;K=F_{2}|D_{2}|^{1/2}$ (16)
where $D_{1}$ ($D_{2}$) is the diagonal matrix of eigenvalues of $Q_{11}$ ($Q_{22}$) associated with an orthonormal matrix of eigenvectors $F_{1}$ ($F_{2}$) sorted such that $\textnormal{sgn}(D_{1})=-\Sigma_{i}$ ($\textnormal{sgn}(D_{2})=\Sigma_{e}$).
In the special case where the system inertia (8) is identity $\Sigma=I$, the conditions (9)-(10) become a linear matrix inequality (LMI) in $Q\succ 0$. In this case, all systems that can become completely symmetric by the transformation (2) are characterized using a convex condition.
In the general case, however, the condition on non-singularity of $Q$ in Theorem 1 is non-convex. Even specifying the system inertia $\Sigma$ (and hence the sign of the eigenvalues of $Q$) a priori does not resolve this issue; except for the positive and negative definite cones, all the other cones in which $Q$ has fixed eigenvalue signs are non-convex. Therefore, it can be difficult to verify the conditions in Theorem 1 numerically in practice.
We obtain necessary conditions for a linear system to be symmetrizable and determine its possible inertia based on a minimal realization in the next section. These conditions are more computationally amenable than Theorem 1.
V Symmetrizability
In this section, we provide necessary conditions for the equations (9) and (10) to have a non-singular solution. First, we use symmetric decomposition to show equation (9) always has a non-singular solution. According to this decomposition, for every square matrix $P$ there are two real symmetric matrices $Q$ and $N$ such that [taussky]
$P=QN,$ (17)
where $Q$ is non-singular. Since $N=Q^{-1}P$ is symmetric, one has $Q^{-1}P=P^{\top}Q^{-1}$, where pre- and post-multiplying $Q$ recovers (9). Therefore, $Q$ is a non-singular solution of (9) if and only if it is the left non-singular factor in a symmetric decomposition of $P$. All these factors can be found analytically using the following proposition.
Proposition 3 ([uh1974])
Let $J=V^{-1}PV=\bigoplus_{i=1}^{b}J_{i}$ be the real Jordan normal form of $P$ and let $E_{i}$ be a matrix of the same size as the Jordan block $J_{i}$ and of the form
$E_{i}=\begin{bmatrix}0&&1\\ &.^{.^{.}}&\\ 1&&0\end{bmatrix}.$ (18)
Then all non-singular left symmetric factors in symmetric decompositions of $P$ are given by
$Q=VU\bigl{(}\bigoplus_{i=1}^{b}\sigma_{i}E_{i}\bigr{)}U^{\top}V^{\top},$ (19)
where $\sigma_{i}\in\{-1,+1\}$ and $U\in\textnormal{Cent}(J)$ is non-singular.
In addition to (9), condition (10) also needs to be satisfied for a system to be symmetrizable, according to Theorem 1. Therefore, we conclude that a system is symmetrizable if and only if its system matrix (7) has a block-diagonal non-singular symmetric left factor as (19), i.e., a factor that satisfies (10). To simplify this condition, we let $P$ have $r:=p+q$ distinct real eigenvalues $\lambda_{1},\,\lambda_{2},\,\dots,\lambda_{p},$ and distinct pairs of complex-conjugate eigenvalues
$(\lambda_{p+1},\bar{\lambda}_{p+1}),\,(\lambda_{p+2},\bar{\lambda}_{p+2}),\,\dots,\,(\lambda_{p+q},\bar{\lambda}_{p+q}),$
regardless of multiplicities. We define
$t_{j}:=\left\{\begin{array}[]{ll}\textnormal{mul}(\lambda_{j}),&\lambda_{j}\in\mathbb{R}\\ 2\,\textnormal{mul}(\lambda_{j}),&\lambda_{j}\not\in\mathbb{R}\end{array}\right.,\quad j=1,2,\dots,r,$
and assume that
$J=V^{-1}PV=\bigoplus_{j=1}^{r}\hat{J}_{j}$
is the real Jordan normal form of $P$, where $\hat{J}_{j}$ contains all the Jordan blocks with eigenvalue $\lambda_{j}$ if $\lambda_{j}\in\mathbb{R}$, and all the Jordan blocks with eigenvalues $\lambda_{j},\bar{\lambda}_{j}$ otherwise. Therefore the blocks with the same real eigenvalues are next to each other and the blocks with the same pairs of complex-conjugate eigenvalues are also adjacent. We also partition $V$ as
$V=\begin{bmatrix}W_{n\times(n+m)}\\ Z_{m\times(n+m)}\end{bmatrix},$ (20)
where $W,Z$ are further partitioned column-wise as follows
$W$ $=\begin{bmatrix}W_{1}&W_{2}&\dots&W_{r}\end{bmatrix},\quad W_{j}\in\mathbb{R}^{n\times t_{j}}$
$Z$ $=\begin{bmatrix}Z_{1}&Z_{2}&\dots&Z_{r}\end{bmatrix},\quad Z_{j}\in\mathbb{R}^{m\times t_{j}}.$ (21)
Theorem 2
If the system represented by (7) is symmetrizable, then
$\textnormal{Rank}(Z\star W)<t_{1}^{2}+\dots+t_{r}^{2}.$ (22)
Proof:
First, we note that the permutation of Jordan blocks does not generate new factors $Q$ in (19). To see this, let $Y=\prod_{k}\mathcal{E}_{k}$ be an arbitrary product of elementary block permutation matrices $\mathcal{E}_{k}$. Then substituting $V$, $U$, $\bigoplus_{i=1}^{b}\sigma_{i}E_{i}$ in (19) by
$VY,\;Y^{-1}UY,\;Y^{-1}\bigl{(}\bigoplus_{i=1}^{b}\sigma_{i}E_{i}\bigr{)}Y$
respectively, does not change the factor $Q$. Therefore, without loss of generality, we assume that the Jordan blocks with the same eigenvalues are placed next to each other in $J$ as follows
$J=\bigoplus_{i=1}^{b}J_{i}=\bigoplus_{j=1}^{r}\hat{J}_{j},$
By invoking Proposition 1, one has
$\textnormal{Cent}(J)=\left\{\bigoplus_{j=1}^{r}\hat{U}_{j}\,\biggl{|}\biggr{.}\,\hat{U}_{j}\in\textnormal{Cent}\bigl{(}\hat{J}_{j}\bigr{)},\;j=1,2,\dots,r\right\}.$
Define
$X:=U\bigl{(}\bigoplus_{i=1}^{b}\sigma_{i}E_{i}\bigr{)}U^{\top}.$ (23)
By regrouping the matrices $\sigma_{i}E_{i}$ as $\bigoplus_{i=1}^{b}\sigma_{i}E_{i}=\bigoplus_{j=1}^{r}\hat{E}_{j}$ where $\hat{E}_{j}$ has the same size as $\hat{J}_{j}$, we may write (23) in the following form
$X$ $=\bigoplus_{j=1}^{r}\hat{U}_{j}\bigl{(}\bigoplus_{j=1}^{r}\hat{E}_{j}\bigr{)}\bigoplus_{j=1}^{r}\hat{U}^{\top}_{j}$
$=\bigoplus_{j=1}^{r}\hat{U}_{j}\hat{E}_{j}\hat{U}^{\top}_{j}$
$:=\bigoplus_{j=1}^{r}X_{j}.$ (24)
Since
$Q=VXV^{\top},$ (25)
[…]
$\left(Z\otimes W\right)\textnormal{vec}(X)=0.$ (26)
By plugging in (V), equation (26) can be written as follows
$\left(Z\star W\right)x=0,$ (27)
where
$\left(Z\star W\right)$ $=\begin{bmatrix}Z_{1}\otimes W_{1}&\dots&Z_{r}\otimes W_{r}\end{bmatrix}$
$x$ $=[\textnormal{vec}(X_{1})^{\top},\dots,\textnormal{vec}(X_{r})^{\top}]^{\top}.$ (28)
However according to (25), $Q$ is non-singular if and only if $X$ is non-singular. Therefore, if equations (9), (10) have a non-singular solution for $Q$, then the Khatri-Rao product (V) has a positive nullity. This condition is equivalent to (22) by the rank-nullity theorem. ∎
Remark 2
The Khatri-Rao product of two matrices is generically full rank [KhatriRank]. Hence according to (22), random linear systems (with system eigenvectors drawn from continuous probability density functions) are neither symmetric nor symmetrizable with probability one.
As varying $L$ does not change the class induced by the equivalence relation $\stackrel{{\scriptstyle\mathclap{\footnotesize\mbox{s}}}}{{\sim}}$, symmetrizability is independent of the realization. Hence, as the null space of the Khatri-Rao product $Z\star W$ encodes all the non-singular solutions $Q$ of the symmetrizability equations (9)-(10) (cf. equation (27)), this space should not change with the realization gains $T$, $K$. To show this fact, let $\hat{P}=LPL^{-1}$ be another realization of the system, where $L$ is a non-singular block matrix as (11). The eigenvectors of $\hat{P}$ are the columns of
$LV=\begin{bmatrix}TW\\ KZ\end{bmatrix}.$
However, one has $(KZ\star TW)=(K\otimes T)(Z\star W)$ where $K\otimes T$ is full rank. Therefore,
$\textnormal{ker}(KZ\star TW)=\textnormal{ker}(Z\star W).$
The next corollary gives more insight in the special case where $P$ has $n+m$ real distinct eigenvalues.
Corollary 1
Assume $P$ has real distinct eigenvalues. The system represented by $P$ is symmetrizable if and only if there is some $x\in\textnormal{ker}(Z\star W)$ with non-zero elements. Furthermore, all the possible system signatures are given by
$i(\Sigma)=\sum_{j=1}^{n+m}\textnormal{sgn}(x_{j}).$
Proof:
When $P$ has real distinct eigenvalues the matrices (18) equal unity, $J$ is diagonal and $\textnormal{Cent}(J)$ is the set of all matrices that are simultaneously diagonalizable with $J$, that is all diagonal matrices. Therefore, all the solutions of (9)-(10) can be written as (25) where $X$ is a non-singular diagonal matrix, whose diagonal elements are given by $x$ in (V). ∎
When $P$ has distinct real eigenvalues, the matrix $Q$ is singular if and only if $x_{j}=0$ holds for some $j=1,2,\dots,m+n$, according to (25). In this case, Corollary 1 makes it easy to isolate the singularities that make the conditions of Theorem 1 non-convex and verify symmetrizability using linear feasibility programs. In particular, the system represented by (7) is symmetrizable with signature $i$, if and only if the following linear program is feasible
$\begin{array}[c]{rl}\text{find}&x\in\mathbb{R}^{n+m}\\ \mbox{subject to}&(Z\star W)\,x=0\\ &e_{j}x_{j}>0,\quad j=1,2,\dots,m+n\end{array}$
for one $e\in\{-1,+1\}^{n+m}$ where $\sum_{j=1}^{n+m}e_{j}=i$.
VI Examples
In this section, we show how symmetrizability extends symmetry in a physical system (Example VI-A), check whether a given system is symmetrizable, determine the possible system signatures, and compute symmetrizing gains (Example VI-B) and extend the analytic solution of an optimal control problem for symmetric systems to symmetrizable systems (Example VI-C).
VI-A A classical multi-tank system
Consider a quadruple-tank process with two inputs (voltages applied to the pumps) and two outputs (levels of the lower tanks) [nunes1998modeling]. This process can be described by the transfer function
$G(s)=\begin{bmatrix}\frac{c_{11}}{1+sT_{1}}&\frac{c_{12}}{(1+sT_{1})(1+sT_{3})}\\ \frac{c_{21}}{(1+sT_{2})(1+sT_{4})}&\frac{c_{22}}{1+sT_{2}}\end{bmatrix},$ (29)
where
$c_{11}$ $=\gamma_{1}k_{1}T_{1}k_{c}/A_{1},$
$c_{12}$ $=(1-\gamma_{2})k_{2}T_{1}k_{c}/A_{1}\neq 0,$
$c_{21}$ $=(1-\gamma_{1})k_{1}T_{2}k_{c}/A_{2}\neq 0,$
$c_{22}$ $=\gamma_{2}k_{2}T_{2}k_{c}/A_{2}.$ (30)
In (29)-(VI-A), $A_{i}$ is the cross-section and $T_{i}$ is the time constant associated with the $i$th tank, constants $k_{i}$ and $\gamma_{i}$ are determined by the $i$th pump and valve settings and $k_{c}>0$. From definition (1), this system is symmetric if and only if
$T_{1}$ $=T_{2},T_{3}=T_{4}\;\textnormal{or}\;T_{1}=T_{4},T_{2}=T_{3}$ (31)
$c_{12}$ =±$=\pm\,c_{21}.$ (32)
However, the first condition (31) is sufficient for the system to be symmetrizable. Hence symmetrizability imposes a less restrictive condition on the tank parameters than symmetry.
VI-B A numerical example
Consider a system described by the minimal state-space equations (6) where
$A$ $=\begin{bmatrix}3.60&-0.63\\ -0.30&3.40\end{bmatrix},$
$B$ $=\begin{bmatrix}-2.14&1.32&3.64\\ -1.86&2.28&9.96\end{bmatrix},$
$C$ $=\begin{bmatrix}0.05&-0.38\\ 2.22&-1.91\\ -0.50&0.47\end{bmatrix},$
$D$ $=\begin{bmatrix}2.20&-1.80&-8.40\\ -2.40&-4.40&-37.20\\ 0.40&1.40&10.20\end{bmatrix}.$
It can be verified that this system is not symmetric, by calculating its transfer function and checking the condition (1). To see whether it is symmetrizable, we calculate the eigenvectors (20) of $P$ in (7). The system matrix $P$ has $n+m=5$ real distinct eigenvalues and
$\textnormal{ker}(Z\star W)=\textnormal{col}\Biggl{(}\begin{bmatrix}1&0&0&0&0\\ 0&1&1&1&1\end{bmatrix}^{\top}\Biggr{)}.$
Therefore, according to Corollary 1, this system is symmetrizable with the signatures $i(\Sigma)\in\{\pm 5,\pm 3\}$. Next, we demonstrate how the symmetrizing gain $K$ is computed to realize a desired signature. Let us choose $i(\Sigma)=-3$ and take a point of the kernel space $\textnormal{ker}(Z\star W)$ with the corresponding sign, e.g.,
$x=\begin{bmatrix}1&-1&-1&-1&-1\end{bmatrix}^{\top}.$
Then $Q$ is calculated as (25) and the gain $K$ is obtained by solving (14), e.g. through (16) as
$K=\begin{bmatrix}2.3460&-0.0576&-2.7399\\ -3.0744&0.0637&-2.1109\\ 0.9380&0.3529&-0.0658\end{bmatrix}.$
One can verify that $H(s)=K^{-1}G(s)K$ is symmetric with $i(\Sigma_{e})=1$.
VI-C An optimal control problem
Consider the linear system described by the minimal state-space equations
$\begin{array}[]{rcl}\dot{x}(t)&=&Ax(t)+Bu(t)+w(t)\\ y(t)&=&Cx(t)+Du(t)\end{array},$ $\quad t\geq 0$ (35)
where $x(0)=0$ and $w(t)\in\mathbb{R}^{n}$ is the disturbance input. We would like to find an output-feedback linear controller that stabilizes the system and minimizes the performance measure
$\mathcal{J}(R,S)=\sup_{w\in\mathcal{W}(S)}\int_{0}^{\infty}y(t)^{\top}Ry(t)+\alpha^{2}u(t)^{\top}Ru(t)dt$
where $\alpha>0$ balances the control effort versus output regulation in the objective function and
[…]
The constants $R,S\succ 0$ are determined based on the dynamics (35) in a way explained shortly. This problem was introduced and solved for relaxation systems in [pates2019optimal], for which $R=I$. A system is of relaxation type, if and only if it admits a completely symmetric realization (one with $\Sigma=I$) and the eigenvalues of $A$ and $-D$ (in any of its minimal realizations) are on the closed left-half plane. We extend this result to symmetrizable systems with $\Sigma=I$ in which the eigenvalues of $A$ and $-D$ are located on the closed left-half plane. These systems include relaxation systems as a proper subset and allow for $R\succ 0$ as any positive definite matrix.
Assume that system (35) is symmetrizable with $\Sigma=I$. Then the LMI (9)-(10) has a solution $Q\succ 0$ which can be used in (13)-(14) to find a similarity transform matrix $T$ and a symmetrizing gain $K$ as follows
$T=Q_{11}^{1/2},\;K=Q_{22}^{1/2}.$
If the eigenvalues of $A$ and $-D$ are located on the closed left-half plane, equations (9)-(10) can be written as
$A^{\top}T^{-2}=T^{-2}A\preceq 0,\;(K^{-1}C)^{\top}=T^{-2}BK,$
$K^{-1}DK\succeq 0,$
which certifies that the symmetrized system (2) with the following state space is of relaxation type
$\dot{\hat{x}}(t)$ $=T^{-1}AT\hat{x}(t)+T^{-1}BKu_{s}(t)+w_{s}(t)$
$y_{s}(t)$ $=K^{-1}CT\hat{x}(t)+K^{-1}DKu_{s}(t),$ (36)
where $y_{s}(t)=K^{-1}y(t)$, $u_{s}(t)=K^{-1}u(t)$, and $w_{s}=T^{-1}w(t)$. Choosing
$R=K^{-2},\;S=T^{-2}$
sets the optimal control problem in the recognized form of [pates2019optimal], in terms of the relaxation system (VI-C), its inputs $u_{s}$, $w_{s}$ and its output $y_{s}$. Based on the results of [pates2019optimal], the performance measure $\mathcal{J}(K^{-2},T^{-2})$ is minimized under the relaxation dynamics (VI-C) with a static output-feedback controller with the closed-form expression
$u_{s}(t)=-\alpha^{-1}(K^{-1}DK-K^{-1}CA^{-1}BK)y_{s}(t).$
Therefore, the optimal feedback controller from $y$ to $u$ is given by
$u(t)=-\alpha^{-1}(D-CA^{-1}B)y(t).$
VII Conclusion
We studied the problem of transforming a linear system into a symmetric system using a static gain. We derived the conditions for this transformation to exist and provided a method to obtain both the symmetrizing gains and symmetric realizations with different signatures. Symmetrizability is a weaker condition than symmetry. Yet, it allows one to apply the simplifying properties of symmetry when controlling symmetrizable systems. To demonstrate this point, we revisited an optimal control problem and extended its analytic solution for symmetric systems to symmetrizable systems. Finally, we showed that linear systems are not symmetrizable generically. Therefore, a future research direction would be to approximate systems by the “nearest” symmetrizable systems.
[…]